# Hype-Adjusted Probability Measure for NLP Stock Return Forecasting

### v2

@@ 3. $3$
$SentAll_{d}=$ $\sum_{i=1}^{k}\mathbbm{1}_{Sent_{d},Sent_{d-1}}\quad\omega_{\text{weight}_{i,d}}SentScore_{i,d}$ (18)
@@
We consider the influence of historical sentiment. The $SentAll_{d}$ variable represents the compound daily sentiment score, taking into account the cumulative influence of historical data. Each term in this equation has the following interpretation:
@@
We hypothesize that when sentiment shifts from positive to negative, the impact becomes more significant as the magnitude of the change increases. This would directly correlate with the weight we aim to introduce here and may warrant further study.
where
  • $\omega_{\text{time memory},d}$ depends on how many days ahead, and whether it is a trading day or not,
  • $\omega_{\text{component weight}}$ depends on the individual market weight of the ticket in its sector.
@@
3.3.1 Continuous Sentiment Equation
To convert the discrete sentiment score equation into a continuous form using integrals, we replace summations with integrals and assume that $Sent_{d}$, $Sent_{d-1}$, and $SentAll_{d}$ represent continuous sentiment functions over time, $t$.
The discrete sum:
$SentAll_{d}=\sum_{i=1}^{k}\mathbbm{1}_{Sent_{d},Sent_{d-1}}\omega_{\text{weight}_{i,d}}SentScore_{i,d}$
can be replaced by an integral form over a continuous time period:
[…]
We now transform each case in the discrete equation to its continuous form:
  1. 1.
    Case 1: When $Sent(t)<0$ and $Sent(t-\Delta t)>0$:
    $SentAll(t)=$ $\,\omega_{\text{today weight}}\cdot Sent(t)$
[…]
  2. 2.
    Case 2: When $Sent(t)>0$ and $Sent(t-\Delta t)<0$:
    $SentAll(t)=$ $\,\omega_{\text{today weight}}\cdot Sent(t)$
[…]
  3. 3.
    Case 3: When $Sent(t)\times Sent(t-\Delta t)\geq 0$:
    $SentAll(t)=$ $\,\omega_{\text{today weight}}\cdot Sent(t)$
[…]
These integrals represent the continuous version of how historical sentiment influences the current sentiment based on the indicator functions $I$, which trigger depending on the changes in sentiment over time. Each term involves weighting past sentiment contributions and the present day’s sentiment to compute the total sentiment.
Note that the period does not need to be daily; the gap interval can be shorter for intraday analysis, such as every hour, allowing for finer granularity in processing and analysis.
@@ 3.4 New NLP Approach and Forecast Results
This section summarizes the machine learning (ML) methodologies, procedures, and forecast results used in the new NLP approach, which is based on the enhanced sentiment score equation.
3.4.1 Methodologies
The primary method introduced in the earlier paper relies on LDA. In this study, however, we use unsupervised ML models as benchmarks to compare the performance of enhanced computational and optimization techniques.
Our objectives are twofold: first, to improve the forecasting accuracy of the original model, and second, to demonstrate the benefits of incorporating arithmetic computations and optimizations beyond standard LDA.
@@ 3.4.1 Methodologies
The optimization algorithms used include models such as LDA, Logistic Regression, Ordinary Least Squares (OLS) regression. The primary focus of this paper is not to fine-tune machine learning models but to demonstrate the improvements gained from the new sentiment score data and procedure.
@@ 3.4.2 Machine Learning Procedures
We employed several machine learning approaches to predict stock market movements based on sentiment data. The sentiment data was first processed by converting the date index into a standardized date format. This data was then merged with stock market data, including daily market returns and volatility metrics.
The methodology involved iterating through various sentiment indicators and performing data splits into training, validation, and test sets. For each indicator, we ran the model up to 1,000 times with different random states to ensure robust results.
Ordinary Least Squares (OLS) regression was used to quantify the relationship between sentiment scores (independent variable $X$) and daily market return or volatility (dependent variable $Y$). The regression provides an R-squared value to measure the proportion of variance in $Y$ explained by $X$. This allows us to evaluate the predictive power of sentiment scores in explaining market dynamics.
Logistic regression was employed to predict the direction of daily market returns (dependent variable $Y$) based on the compound sentiment score (independent variable $X$). Here, $Y$ was a binary variable indicating positive or negative market movements. The accuracy of the logistic regression model served as a key metric for assessing its performance.
The optimal models were selected based on the highest validation accuracy (for logistic regression) and R-squared values (for OLS regression) across iterations. These models were subsequently evaluated on the test dataset, with metrics including accuracy, precision, recall, and confusion matrices. The results highlighted the most effective sentiment indicators for predicting market movements, offering insights into the relationship between sentiment data and market behavior.
Sample table showing modified sentiment scores based on different parameters (for equation 18) from the new equation:
@@
The model results are analyzed in Section 5.1.
4 Hype-Adjusted Probability Measure
4.1 Intuition of the New Measure
While we previously proposed a sentiment equation in discrete form, we now extend it to a continuous framework to capture market dynamics with greater precision. The limitations of discrete methods prompt us to provide mathematical justifications alongside market implications, addressing parameters like the “3 weights" introduced earlier.
In the context of financial markets, hype refers to the amplification of attention or sentiment around a particular stock, sector, or event that exceeds its fundamental or intrinsic importance. Hype is often fueled by disproportionate media coverage, speculative behavior, or investor overreaction to news.
@@ 4.1 Intuition of the New Measure
    Market Reactions: Corresponding anomalies in price movement and volatility, such as sharp increases or dramatic swings, that deviate from historical patterns.
@@
    Imbalance in Representation: Evidence of over or under representation in news coverage compared to the stock’s weight in its sector (e.g., market capitalization).
@@
Recall the “3 weights":
@@
Also, recall that we have shown from figures that a slight negative skew of market news sentiments is observed and there exists an imbalance of news report counts vs. component weights of Nvidia among the semiconductor sector tickers.
In addition, the news source bias weights are currently being investigated for further research. For example, in America, major news media have political leanings, CNN is more left-wing and promotes clean energy and gun control, while Republican-favored Fox argues for the opposite tone. We wish to quantify the bias instead of relying solely on polls and surveys.
@@
Notice, while we do not assert a direct positive correlation between news count increases and price or volatility changes, these observations suggest that market hype can significantly impact these variables, which inspires the proposal of the hype-adjusted probability measure, $\mathbb{P}^{a}$.
4.2 Hype-Adjusted Probability Theory
@@
Based on the intuitions, we define the hype-adjusted probability measure, $\mathbb{P}^{a}$, to account for market sentiment and the role of news weight biases.
We choose volatility as the target of prediction, which aligns with the observation results from sections 3.4 and 4.1. We also discuss the relationship between the hype-adjusted probability measure and the volatility and how to apply this for market price forecasting.
Definition: Hype-adjusted Probability Measure
The hype-adjusted probability measure, denoted $\mathbb{P}^{a}$, is a probability measure that incorporates sentiment biases by adjusting the physical measure $\mathbb{P}$. The probability space for hype-adjusted probability measure is a triplet $(\Omega,\mathcal{F},\mathbb{P}^{a})$. Under $\mathbb{P}^{a}$, the effects of news weight, bias from media sources, and sentiment shifts are inherently accounted for, allowing a more accurate representation of expected market behaviors influenced by hype.
4.2.1 Change from Physical to Hype-Adjusted Probability Measure
We get the inspiration of change of measure from the physical measure to Harrison and Kreps’ risk-adjusted measure (1979) and to Geman’s Forward measure (1989) as introduced in the literature review section 2.2.
For simplicity, we illustrate the change of measure by considering only the "Capital vs. News Counts" factor of the “3 weights."
  • Let $L_{Cap}(\omega)$ denote the capital weight of a given event (ticker).
  • Let $L_{News}(\omega)$ denote the news weight of a given ticker.
As shown in Appendix A, it is clear that:
$\sum L_{Cap}(\omega)=\sum L_{News}(\omega)=1$
Hence, we assign
$\mathbb{P}(\omega)=L_{News}(\omega)\quad\text{and}\quad\mathbb{P}^{a}(\omega)=L_{Cap}(\omega),$
which leads to:
$\sum\mathbb{P}(\omega)=\sum\mathbb{P}^{a}(\omega)=1.$
It follows almost surely that no weight is equal to zero. In other words:
$\text{Prob}(L_{Cap}(\omega)=0)=0\quad\text{and}\quad\text{Prob}(L_{News}(\omega)=0)=0.$
Therefore, by the Equivalent Measure Theorem, the two probability measures agree on which sets in $\mathbb{F}$ have probability zero. Intuitively, an neutral sentiment score ($=0$) under $\mathbb{P}$ is still neutral under $\mathbb{P}^{a}$.
The change of measure theorem states that, since $\mathbb{P}$ and $\mathbb{P}^{a}$ are equivalent, the Radon-Nikodym derivative
$\mathbb{Z}(\omega)=\frac{d\mathbb{P}(\omega)}{d\mathbb{P}^{a}(\omega)}$ (20)
exists. For any $X\in\mathcal{F}$, we have:
@@ 4.2.1 Change from Physical to Hype-Adjusted Probability Measure
This adjustment allows us to account for market hype by converting the original news weight of a particular ticker to its correct proportion within the sector.
Remark - Sentiment Adjustment:
For the hype-adjusted probability measure $\mathbb{P}^{a}$, the Radon-Nikodym derivative $\mathbb{Z}(\omega)$ adjusts the impact of news on tickers based on their level of media coverage. Specifically:
  • For over-hyped tickers (e.g., NVDA.QQ), the news count impact is discounted, represented by $\mathbb{Z}(\omega)<1$.
  • For under-represented tickers (e.g., MPWR.OQ), the news count impact is amplified, represented by $\mathbb{Z}(\omega)>1$.
This adjustment balances sentiment influence across both highly publicized and lesser-known tickers.
Recall, under $\mathbb{P}$, we have equation 19 as :
$\text{Sentiment}(t)$ $=SentAll(t)$ (21)
[…]
Under $\mathbb{P}^{a}$, we can rewrite equation 19 as:
[…]
where $\hat{\text{Sentiment}}(t)$ represents the adjusted sentiment score of a given period under $\mathbb{P}^{a}$.
We then obtain the expected sentiment score $\hat{\text{Sentiment}}(t)$ over time $t$ as:
[…]
where:
  • $\frac{d\mathbb{P}}{d\mathbb{P}^{a}}$: the Radon-Nikodym derivative, representing the adjustment from the original measure $\mathbb{P}$ to the hype-adjusted measure $\mathbb{P}^{a}$.
  • $\mathbbm{1}(Sent(s),Sent(s-\Delta s))$: an indicator function that activates based on sentiment shifts over the interval $[s,s-\Delta s]$.
  • $\omega_{\text{weight}}(s)$: the weighting function for sentiment impact at time $s$.
  • $SentScore(s)$: the sentiment score at time $s$.
  • $\mathcal{F}_{t}$: the filtration up to time $t$, representing the information available in the market.
Thus, the sentiment score $\hat{\text{Sentiment}}(t)$ under the hype-adjusted measure $\mathbb{P}^{a}$ is defined as:
Theorem 4.1 (Conditional Expectation of Sentiment).
The Conditional Expectation of Sentiment at time T under the hype-adjusted probability measure is:
[…]
@@
where $\hat{SentScore}(T)$ represents the adjusted sentiment score under $\mathbb{P}^{a}$.
In this form, the adjusted measure $\mathbb{P}^{a}$ captures market sentiment dynamics directly through the sentiment score $\text{SentScore}(s)$ alone. The absence of explicit weighting and indicator functions in Equation 26 highlights that $\mathbb{P}^{a}$ automatically adjusts for sentiment influence and bias over time, streamlining the overall computation.
We propose:
Theorem 4.2 ($\mathbb{P}^{a}$ martingale).
There exits a $\mathbb{P}^{a}$ martingale, such that:
$\mathbb{P}^{a}\sim\mathbb{P}$
and
$\mathbb{E}^{\mathbb{P}^{a}}\left[\frac{\partial\hat{\text{Sentiment}}(s)}{\partial\text{Volatility(s)}}\mid\mathcal{F}_{t}\right]=\frac{\partial\hat{\text{Sentiment}}(t)}{\partial\text{Volatility(t)}},\quad\forall s>t,$ (27)
where $\sigma(s)$ is the volatility of underlying security prices.
4.2.2 Hype-Adjusted Measure Under Discrete Binomial Tree Model
The discrete binomial tree model provides a clear framework to illustrate the concept of changing measures in a discrete setting. Traditionally, the binomial tree represents potential future movements of asset prices or volatility over a series of discrete time intervals. Here, we adapt this model to transition from the physical measure $\mathbb{P}$ to the hype-adjusted measure $\mathbb{P}^{a}$ by adjusting the probabilities associated with upward and downward volatility movements, incorporating the "Capital vs. News Counts" factor as weights.
Unlike a standard binomial asset pricing tree model, we have:
  • At each time step, the volatility can move up by a factor $u$ with probability $p$ or down by a factor $d$ with probability $1-p$.
  • The volatility at time $t+\Delta t$ is given by:
    $\sigma_{t+\Delta t}=\begin{cases}\sigma_{t}u&\text{with probability }p,\\ \sigma_{t}d&\text{with probability }1-p.\end{cases}$
To introduce the hype-adjusted measure, define:
  • $L_{Cap}$: the capital weight associated with each event (ticker).
  • $L_{News}$: the news weight assigned to each event (ticker).
Under the physical measure $\mathbb{P}$, the probabilities for an upward and downward movement are given by $p=L_{News}(u)$ and $1-p=L_{News}(d)$, respectively. The hype-adjusted measure $\mathbb{P}^{a}$ adjusts these probabilities as follows:
$p=L_{Cap}(u)\quad\text{and}\quad 1-p=L_{Cap}(d).$
This adjustment captures the impact of sentiment-driven volatility changes, with $L_{Cap}$ and $L_{News}$ modulating the probabilities under the hype-adjusted measure.
The Radon-Nikodym derivative $\mathbb{Z}(\omega)$ allows us to transition from $\mathbb{P}$ to $\mathbb{P}^{a}$. In the binomial tree context, this derivative can be expressed at each node as:
$\mathbb{Z}(u)=\frac{L_{Cap}(u)}{L_{News}(u)},\quad\mathbb{Z}(d)=\frac{L_{Cap}(d)}{L_{News}(d)}.$
These adjusted probabilities reflect how market sentiment biases (e.g., disproportionate media focus on certain tickers) are incorporated into the measure.
For volatility $\sigma_{t}$, the expected value at the next time step under $\mathbb{P}^{a}$ is:
$\mathbb{E}^{\mathbb{P}^{a}}[\sigma_{t+\Delta t}\mid\sigma_{t}]=\sigma_{t}\left(L_{Cap}(u)u+L_{Cap}(d)d\right).$
In contrast, under $\mathbb{P}$, it would be:
$\mathbb{E}^{\mathbb{P}}[\sigma_{t+\Delta t}\mid\sigma_{t}]=\sigma_{t}\left(L_{News}(u)u+L_{News}(d)d\right).$
This adjustment aligns the asset’s future volatility expectations with the corrected market sentiment weightings, effectively capturing hype adjustments in a discrete framework.
4.2.3 Sentiment-Driven Volatility and Hype-Adjusted Pricing
Building upon the discrete analysis, our work leverages established relationships between volatility, stock prices, and sentiment, as demonstrated in models such as the Black-Scholes-Merton (BSM) model. In the BSM framework, volatility is a critical parameter that directly influences option pricing. Traditionally, volatility is viewed as a measure of market risk and is calculated based on historical price movements. However, we have shown that sentiment shifts, particularly those driven by hype or disproportionate news coverage, can significantly impact volatility.
Theorem 4.3 (Sentiment-Driven Volatility Fluctuations).
Let $\sigma(t)$ denote the volatility of an underlying asset and $\text{Sentiment}(t)$ represent the sentiment score at time $t$. We assume that fluctuations in volatility are sentiment-driven. Empirical analysis suggests a correlation between changes in sentiment $d\text{Sentiment}(t)$ and adjustments in volatility $d\sigma(t)$, such that:
$d\sigma(t)\approx\alpha\,d\text{Sentiment}(t),$ (28)
where $\alpha$ is a sensitivity parameter derived from historical relationships between sentiment and volatility, as proposed in equation 27.
This extension allows for sentiment-adjusted volatility $\hat{\sigma}(t)$, defined as:
$\hat{\sigma}(t)=\sigma(t)+\alpha\,d\text{Sentiment}(t),$ (29)
which reflects sentiment-driven adjustments.
The sentiment-adjusted volatility $\hat{\sigma}(t)$ serves as a dynamic input that adapts to sentiment variations, providing a refined approach to volatility forecasting in sentiment-sensitive markets.
Theorem 4.4 (Sentiment-Adjusted Stochastic Price Dynamics).
Under the hype-adjusted probability measure $\mathbb{P}^{a}$, the distribution of price trajectories incorporates sentiment-driven adjustments. The stochastic differential equation governing the asset price $S_{t}$ is modified to reflect sentiment-adjusted volatility, as follows:
$\frac{dS_{t}}{S_{t}}=r\,dt+\hat{\sigma}(t)\,dW_{t}^{\mathbb{P}^{a}},$ (30)
where $dW_{t}^{\mathbb{P}^{a}}$ represents the Wiener process under the hype-adjusted measure $\mathbb{P}^{a}$, and $\hat{\sigma}(t)$ is the volatility adjusted for sentiment effects.
This modification results in a set of stochastic trajectories that align price forecasts with sentiment-influenced market expectations, enabling a more accurate representation of market dynamics in sentiment-sensitive contexts.
By integrating the hype-adjusted measure $\mathbb{P}^{a}$ and sentiment-sensitive adjustments to volatility, we enhance traditional pricing models to better reflect the influence of market hype. This framework not only updates the pricing formula but also provides a probabilistic model for sentiment-adjusted price paths. Such insights enable more accurate derivative pricing in sentiment-driven markets and support refined investment strategies during periods of heightened market hype.
For example, with $\sigma(t)$ now incorporating sentiment adjustments in equation 29, we can update the standard Black-Scholes pricing formula for a European call option $C(S,t)$. The sentiment-adjusted volatility $\hat{\sigma}(t)$ replaces the traditional volatility $\sigma$ in the BSM formula, producing a sentiment-enhanced price:
$C(S,t)=SN(d_{1})-Ke^{-r(T-t)}N(d_{2}),$
where:
$d_{1}$ $=\frac{\ln\left(\frac{S}{K}\right)+\left(r+\frac{{\hat{\sigma}(t)}^{2}}{2}\right)(T-t)}{\hat{\sigma}(t)\sqrt{T-t}},$
$d_{2}$ $=d_{1}-\hat{\sigma}(t)\sqrt{T-t},$
At time $t$, we have information granted by the filtration $F_{t}$. We may decide which volatility $\sigma$ model or method to obtain the stock price $S(T)$ at future time $T$ and compare the results.
[…]
For example, at time $t$, we can derive volatility using one of the following methods:
  • The implied volatility $\sigma_{\text{Imp}}$ from the Black-Scholes equation.
  • The adjusted volatility $\hat{\sigma}$, incorporating news-based sentiment adjustments.
These choices lead to various trading strategies, such as a variance swap.
In summary, this section defines $\mathbb{P}^{a}$, from the change of measure from $\mathbb{P}$ to the analysis under the discrete binomial model, and the discussion for sentiment-driven volatility and hype-adjusted pricing.
5 Results and Future Work
5.1 Model Results
We observed that the expected average difference between the original volatility and the adjusted one, based on the weights provided in the dataset, is approximately $-0.0068$ (or $-0.68\%$). Significant improvements were observed with the new NLP approach utilizing the enhanced sentiment scoring equation.

### v4

@@ 3.3 Adjusted Sentiment Score Equation
$SentAll_{d}=$ $\sum_{i=1}^{n}\mathbbm{1}_{Sent_{d},Sent_{d-1}}\quad\omega_{\text{weight}_{i,d}}SentScore_{i,d}$ (13)
@@
We consider the influence of historical sentiment. The $SentAll_{d}$ variable represents the compound daily sentiment score of $n$ tickers, taking into account the cumulative influence of historical data.
@@
We hypothesize that when sentiment shifts from positive to negative, the impact becomes more significant as the magnitude of the change increases. This would directly correlate with the weight we aim to introduce here.
@@ 3.4 New NLP Approach and Forecast Results
This section summarizes the ML methodologies, procedures, and forecast results used in the new NLP approach, which is based on the enhanced sentiment score equation.
3.4.1 Methodology
The primary method introduced in Deveikyte et al relied on Latent Dirichlet Allocation [deveikyte2022]. In contrast, we use in this study unsupervised ML models as benchmarks to compare the performance of enhanced computational and optimization techniques.
Our objectives are twofold: first, improve the forecasting accuracy of the original model, and second, demonstrate the benefits of incorporating arithmetic computations and optimizations beyond standard LDA.
@@ 3.4.1 Methodology
The optimization algorithms used include models such as LDA, Logistic Regression, Ordinary Least Squares (OLS) regression. The primary focus of this paper is not to fine-tune ML models but to demonstrate the improvements gained from the new sentiment score data and procedure.
@@ 3.4.2 Machine Learning Procedures
We employ several ML approaches to predict stock market movements based on sentiment data. The sentiment data is first processed by converting the date index into a standardized date format. This data is then merged with stock market data, including daily market returns and volatility metrics.
The methodology involves iterating through various sentiment indicators and performing data splits into training, validation, and test sets. For each indicator, we run the model up to 1,000 times with different random states to ensure robust results.
Ordinary Least Squares (OLS) regression is used to quantify the relationship between sentiment scores (independent variable $X$) and daily market return or volatility (dependent variable $Y$). The regression provides an R-squared value to measure the proportion of variance in $Y$ explained by $X$. This allows us to evaluate the predictive power of sentiment scores in explaining market dynamics.
Logistic regression is employed to predict the direction of daily market returns (dependent variable $Y$) based on the compound sentiment score (independent variable $X$). Here, $Y$ is a binary variable indicating positive or negative market movements. The accuracy of the logistic regression model serves as a key metric for assessing its performance.
The optimal models are selected based on the highest validation accuracy (for logistic regression) and R-squared values (for OLS regression) across iterations. These models are subsequently evaluated on the test dataset, with metrics including accuracy, precision, recall, and confusion matrices. The results highlight the most effective sentiment indicators for predicting market movements, offering insights into the relationship between sentiment data and market behavior.
Table 4 shows a sample of modified sentiment scores based on different parameters in equation 13.
@@
The model results are analyzed in Section 5.
4 A Hype-Adjusted Probability Measure
In the context of financial markets, hype refers to the amplification of attention or sentiment around a particular stock, sector, or market event that exceeds its fundamental or intrinsic importance. Hype is often fueled by disproportionate media coverage, speculative behavior, or investor overreaction to news.
@@ 4 A Hype-Adjusted Probability Measure
    Market Over Reactions: Corresponding anomalies in price movement and volatility, such as sharp increases or dramatic swings, that deviate from historical patterns.
@@
    Imbalance in Representation: Evidence of over or under-representation in news coverage compared to the stock’s weight in its sector (e.g., market capitalization).
@@
4.1 Intuition of the New Measure
Recall the 3 criteria:
@@ 4.1 Intuition of the New Measure
Also, recall that we have shown from figure 1 that a slight negative skew of market news sentiments is observed and there exists an imbalance of news report counts vs. component weights of Nvidia among the semiconductor sector tickers.
In addition, the news source bias weights are currently being investigated for further research. For example, in America, major news media have political leanings, CNN is more left-wing and promotes clean energy and green targets, while Republican-favored Fox argues for the opposite tone. We wish to quantify the bias instead of relying solely on polls and surveys.
@@
Notice, while we do not assert a direct positive correlation between news count increases and price or volatility changes, these observations suggest that market hype can significantly impact these variables, which inspires the proposal of a hype-adjusted probability measure, $\mathbb{P}^{a}$.
4.2 Construction of A Hype-Adjusted Probability Measure
@@
Based on the previous intuitions, we define a hype-adjusted probability measure, $\mathbb{P}^{a}$, to account for market sentiment and the existence of news biases.
Hype-adjusted Probability Measure:
We consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$ where $\Omega$ is the set of states of nature $\omega$, $\mathcal{F}$ is the filtration of information, and $\mathbb{P}$ the physical probability measure. We define a new probability measure $\mathbb{P}^{a}$ on $(\Omega,\mathcal{F})$ by assigning new weights to the states of nature, where the effects of news weight and bias from media sources are corrected, reducing the over / under representation in the news of some particular stocks of a sector.
We get the economic inspiration of change of measure from the physical measure $\mathbb{P}$ to a Harrison and Kreps (1979) “risk-neutral" / equivalent-martingale measure; and Geman’s Forward measure (1989). Note that in these two famous cases like in our setting below, the change of measure and its weights reflect the economic problem to address.
We suppose that there are 3 states of nature $\omega_{1},\omega_{2},\omega_{3}$ in the universe $\Omega$, labeled as Up, Level, and Down.
We know from measure theory that the relationship between a probability measure $\mathbb{P}$ on $(\Omega,\mathcal{F})$ and an equivalent probability measure $\mathbb{P}^{a}$ has the following form:
$Z=\frac{d\mathbb{P}^{a}}{d\mathbb{P}},$ (14)
$Z(\omega)=\frac{\mathbb{P}^{a}(\omega)}{\mathbb{P}(\omega)}.$ (15)
where $Z$ is a random variable with
$\mathbb{E}_{\mathbb{P}}[Z]=1.$ (16)
In order to define $Z$, we need to choose $Z(\omega_{1})$, $Z(\omega_{2})$, and $Z(\omega_{3})$.
We consider now two companies, Nvidia and Intel, and their respective stock and sentiment performances:
  • Nvidia: Today, the stock increased by 10%, but the sentiment increased by 20%. Nvidia is thus considered “overhyped", we choose:
    $Z(\omega_{1})<1$
  • Intel: The stock decreased by 20%, but the sentiment only decreased by 10%. Intel is thus considered “underhyped", we choose:
    $Z(\omega_{3})>1.$
1. Up State ($w_{1}$):
For Nvidia, which is overhyped, we choose $Z(\omega_{\text{up}})<1$, reducing the probability of the Up state. This reflects the fact that the market sentiment is too optimistic relative to the actual stock performance.
$\mathbb{P}^{a}(\omega_{\text{up}})=Z(\omega_{\text{up}})\cdot\mathbb{P}(\omega_{\text{up}}),$
such that
$\mathbb{P}^{a}(\omega_{\text{up}})<\mathbb{P}(\omega_{\text{up}}).$
2. Down State ($w_{3}$):
For Intel, which is underhyped, we choose $Z(\omega_{\text{down}})>1$, increasing the probability of the Down state to reflect the understated sentiment relative to the actual performance:
$\mathbb{P}^{a}(\omega_{\text{down}})=Z(\omega_{\text{down}})\cdot\mathbb{P}(\omega_{\text{down}}),$
such that
$\mathbb{P}^{a}(\omega_{\text{down}})>\mathbb{P}(\omega_{\text{down}}).$
3. Level State ($w_{2}$):
In the “Level" state, sentiment and stock price changes are neutral, and no adjustment is needed. The value we assign to $Z(\omega_{\text{Level}})$ is such that
$\mathbb{E}_{\mathbb{P}}[Z]=1.$
The same construction can be extended to include more stocks.
Recalling its definition given in equation 4, the conditional expectation under the hype-adjusted probability measure can be expressed in integral form as:
@@ 3. Level State ($w_{2}$):
where $X(\omega)$ represents the sentiment score for each state $\omega$, and the expectation is taken under the adjusted measure $\mathbb{P}^{a}$.
Additionally, the expectation of $X$ under $\mathbb{P}^{a}$ can also be written in terms of the Radon-Nikodym derivative $Z$:
$\tilde{\mathbb{E}}[X]=\mathbb{E}[X\cdot Z],$ (18)
where $\mathbb{E}[X\cdot Z]$ is the expectation under the original measure $\mathbb{P}$, weighted by $Z$.
Finally, the adjusted sentiment score $X$ under $\mathbb{P}^{a}$ can be rewritten in terms of its relationship with $Z$, with $s<t$, as:
$\mathbb{E}^{\mathbb{P}^{a}}[X\mid\mathcal{F}(s)]=\frac{1}{Z(s)}\mathbb{E}[X\cdot Z(t)\mid\mathcal{F}(s)].$ (19)
This adjustment allows us to account for market hype by converting the original news weight of a particular ticker to its correct proportion within the sector. The following section 5 presents the improved NLP forecasting results using a hype-adjusted probability measure.
5 Discussion
The results presented in this section highlight the effectiveness of the proposed sentiment-based framework in improving market return and volatility forecasting. By integrating advanced sentiment modeling techniques, we observe substantial accuracy improvements, which pave the way for innovative theoretical contributions and practical applications in financial analysis.
5.1 Results
Under a hype-adjusted probability measure of the selected date range, we calculate that the expected average difference between the original volatility and the adjusted one, based on the weights provided in the semiconductor dataset, is approximately $-0.0068$ (or $-0.68\%$). We note however, that despite this small number, the accuracy of the prediction was improved by $8\%$.